OpenStudy (anonymous):

What is the simplified form of: 8x4/12y3 . 20y8/6x6

OpenStudy (anonymous):

help please and fast:(

OpenStudy (ghazi):

\[\frac{ 8x^4 }{ 12 y^3 }*\frac{ 20y^8 }{ 6x^6}={ 20 x^{-4} }*y^2/9\]

OpenStudy (anonymous):

huh

OpenStudy (ghazi):

what?

OpenStudy (ghazi):

just use \[\frac{ x^a }{ x^b }= x^{a-b}\]

OpenStudy (anonymous):

simpler please

OpenStudy (ghazi):

well it is the property ...you have to subtract the power if bases are same